OpenStudy (anonymous):

I NEED HELP PLEASE!!! Simplify: -4b(5a – 2b ) + 4(-3ab – 5b2)

OpenStudy (waheguru):

-32ab-20b2-8b^2

OpenStudy (barrycarter):

http://www.wolframalpha.com/input/?i=-4b%285a+%E2%80%93+2b+%29+%2B+4%28-3ab+%E2%80%93+5b^2%29 under "alternate forms", assuming that last 2 means "squared"

OpenStudy (anonymous):

-20ab+8b^2-12ab-20b^2 =-12b^2-32ab

OpenStudy (waheguru):

ajatri ris correct if the 2 was a square

OpenStudy (anonymous):

yes the 2 is ^2

OpenStudy (anonymous):

thank you guys can anyone help me on this problem ? Simplify: 3a – {5b – 2[ b – ( 1- 2b)]}

OpenStudy (anonymous):

3a+b-2

OpenStudy (paxpolaris):

\[-4b(5a – 2b ) + 4(-3ab – 5b^2)= -4b(5a – 2b ) - 4b(3ab + 5b)\]\[=-4b \left( 5a+3a+5b-2b \right)=-4b(8a+3b)= \Large -32ab-12b^2\]